# Neutron Star with Dark Matter Admixture: A Candidate for Bridging the Mass Gap

### v1

@@ Abstract
Concerning the stability condition, we employ the method developed by Henriques, Liddle and Moorhouse which used for the study of boson-fermion star [63, 64, 65]. This method has been elaborated and extended through the years for similar studies [66, 68, 69, 67]. According to this method, the stability analysis is carried out by examining the behaviour of baryons and DM particles, yet fixing the total mass $M$ values. In particular, the stability curve is formed with the pair of central values of pressures $\{P_{c}^{\rm NS},P_{c}^{\rm DM}$} exactly in the point where the number of particles reached the minimum and maximum values. This critical curves identify the transition from linear, stable and unstable with respect to perturbations that conserve the mass and the particle number. Hence fulfill the following conditions [67]:
@@
where $N_{b}$ and $N_{\chi}$ are the numbers of baryons and DM particles respectively. The stability analysis can be summarized as follows [67]: Sequences of stable equilibrium configurations that start from purely DM star have the feature that the number of DM particle $N_{\chi}$ first decreases (as a function of the central pressures $P_{c}^{\rm DM}$ or $P_{c}^{\rm NS}$) up to the critical point and the number of baryons $N_{b}$ increases. In general, equilibrium sequences can consist of continuous regions of stability and instability. In this case, therefore, we can have more than one branch of stability.
@@
Figure 1: Equilibrium configurations of NS and DM admixture, as a function of the central pressures $P_{\rm DM}^{c}$ and $P_{\rm NS}^{c}$, with interaction strength $y=0.05$ MeV-1 and DM mass $m_{\chi}=1500$ MeV. The colored lines correspond to models with the same total mass. The black solid line depicts the boundary between stable and unstable region. The inner figure visualizes more clearly the stable region for very low values of the central pressures.
@@
To be more specific, in Fig. 1, we present the equilibrium configurations of a NS with DM admixture and an interaction strength of $y=0.05$ MeV-1 and a DM mass of $m_{\chi}=1500$ MeV. The various colored lines represent models with the same total mass, while the black solid line marks the boundary between the stable and unstable regions. The black dot and the red dot correspond to the purely neutron matter and purely DM star cases (with the corresponding masses). This figure clearly demonstrates that the right combination of NS matter and DM can result in stable compact objects that span the mass-gap range. Another clear conclusion is that, within the aforementioned region, the larger the mass of the star, the higher the proportion of DM it contains. This is expected because, while the equation of state limits the maximum mass of a NS, no such restriction applies to a DM star. Depending on its mass and interactions, a DM star’s mass can grow significantly larger.
@@
However, since the radius of these stars is not easily detectable due to the formation of the DM halo, we focus on the tidal deformability of the star: $\Lambda=\frac{2}{3}k_{2}\left(\frac{c^{2}R}{GM}\right)^{5}$, where $k_{2}$ is the tidal Love number [70]. Thus, in Fig. 3, we plot the dimensional $\Lambda$ for three different masses. Obviously, the values it takes are significantly larger compared to those of NSs. Since $\Lambda$ is a quantity that can be determined from observations, it serves as an important tool for detecting these objects. It is important to note that, for a given mass, the range of $\Lambda$ values can span 2-3 orders of magnitude. This outcome arises from the fact that these objects, in their stable state, can vary significantly in size, particularly in terms of their radius.
@@
Finally, we try to shed light on two critical issues concerning the creation and detection of these objects.When and how could supramassive compact objects made of exotic fermions have formed? A relevant discussion is provided in Ref. [71]. For example, one possibility involves the accretion of DM into NS through the primordial formation of DM clumps surrounded by ordinary matter. Another possibility is the formation of dark compact objects originating from DM perturbations that grow from primordial over-densities.
Finally, another possibility is copious production and capture of DM in the core-collapse supernova of NSs progenitor [72]. In any case, there is no strong theoretical prediction that firmly argues against the creation of these hybrid NS-DM objects. The last proposition leads to the most critical question: how could we locate these objects? The most powerful tool in this case is gravitational lensing, which relies on the space-time distortion caused by these objects. Another possibility is detecting a potential merger event through the well-known detectors of the LIGO-Virgo-KAGRA Collaboration.
This includes two scenarios: a) the merger between two dark compact objects, or b) the merger between one of these dark compact objects and another compact object, such as a NS or a BH. The corresponding gravitational wave signals will carry valuable information about the structure of these objects, providing insights into their composition and internal dynamics [73].
@@
Summarizing this study, we propose an alternative way to explain the possible existence of compact objects in the mass-gap region. We consider that there may be hybrid states of coexisting DM and hadronic matter in the form of stable compact stars in the universe, which, among other things, cover the mass-gap region. The formation of these objects allows for multiple interpretations, and despite its significance, it lies outside the scope of this study.
@@
Acknowledgments — This work was supported by the Czech Science Foundation (GACR Contract No. 21-24281S).

### v2

@@ 2.5 Stability Conditions
In this study, we employ an alternative method to investigate the stability of these hybrid objects. Specifically, we use the method developed by Henriques, Liddle, and Moorhouse, which was applied to the study of boson-fermion stars [95, 96, 97]. This method has been developed and refined over the years for similar studies [98, 100, 101, 99, 91]. Using this method, the stability analysis is performed by examining the behavior of baryons and dark DM particles, while keeping the total mass $M$ fixed. Specifically, the stability curve is constructed using the pair of central pressure values $\{P_{c}^{\rm NS},P_{c}^{\rm DM}$} corresponding to the points where the number of particles reaches its minimum and maximum values. These critical curves identify the transition between linear, stable, and unstable states with respect to perturbations that conserve both mass and particle number. Therefore, they satisfy the following conditions [99]
@@
where $N_{b}$ and $N_{\chi}$ represent the numbers of baryons and DM particles, respectively. These numbers are given by the expression [102]
$N_{i}=4\pi\int_{0}^{R_{i}}n_{i}(r)\frac{r^{2}dr}{\sqrt{1-2GM(r)/rc^{2}}},\quad i=b,\chi$ (19)
where $R_{i}$ is the radius of the corresponding fluid.
The stability analysis can be summarized as follows [99, 91]: Sequences of stable equilibrium configurations starting from a purely DM star exhibit the characteristic that the number of DM particles $N_{\chi}$ initially decreases (as a function of the central pressures $P_{c}^{\rm DM}$ or $P_{c}^{\rm NS}$) until reaching the critical point, while the number of baryons $N_{b}$ increases. In general, equilibrium sequences may include continuous regions of both stability and instability. Therefore, in this case, multiple branches of stability can exist.
3 Results and Discussion
@@
Figure 1: Equilibrium configurations of NS and DM admixture, as a function of the central pressures $P_{\rm DM}^{c}$ and $P_{\rm NS}^{c}$, with interaction strength $y=0.05$ MeV-1 and DM mass $m_{\chi}=1500$ MeV. The colored lines represent models with the same total mass, while the black solid line indicates the boundary between the stable and unstable regions. The inner figure visualizes more clearly the stable region for very low values of the central pressures.
Figure 2: The figure depicts the mass for the DM component (solid lines), separately with the NS component (dashed lines), of the compact object as a function of the central pressure of the NS matter $P_{\rm NS}^{c}$, for a three fixed values of the hybrid compact object.
Figure 3: The dependence of the number of DM particles $N_{\rm DM}$ and baryons $N_{\rm NS}$ on the pressure $P_{\rm NS}^{c}$ for equilibrium configurations of equal mass: a) $M=3.12\ M_{\odot}$, b) $M=4.18\ M_{\odot}$, c) $M=5.24\ M_{\odot}$. The solid lines indicate the stable branches, while the dashed lines are the unstable ones. The x appearing on each curve denotes the last stable configuration (the corresponding central pressures, in ${\rm MeV/fm}^{3}$, are also indicated in each case).
@@
To be more specific, in Fig. 1, we present the equilibrium configurations of a NS with DM admixture and an interaction strength of $y=0.05$ MeV-1 and a DM mass of $m_{\chi}=1500$ MeV. The different colored lines represent models with the same total mass, while the black solid line delineates the boundary between the stable and unstable regions. The black and red dots correspond to the cases of pure neutron star and pure DM star, respectively, along with their corresponding masses. This figure clearly demonstrates that the right combination of NS matter and DM can result in stable compact objects that could exist inside the mass-gap range.
In Fig. 2, we illustrate the relationship between the individual mass of a NS and DM for three cases of hybrid objects, plotted as a function of central pressure of the NS matter $P_{\rm NS}^{c}$. These cases represent stable states corresponding to fixed total masses of $3.12$, $4.18$, and $5.24\ M_{\odot}$. As expected, for small values of the hadronic pressure (and correspondingly high values of DM pressure) at the center of the hybrid object, the proportion of DM is overwhelmingly greater compared to that of the NS. However, as the hadronic pressure increases, the proportion of DM decreases, and the NS of the hybrid object approaches its known dimensions in terms of mass and radius. It is worth emphasizing that all these configurations describe stable states. Obviously, these results are sensitive to the fundamental parameters of the problem, which are the mass of the DM and the strength of its interaction.
Another clear conclusion is that, within the aforementioned region, the larger the mass of the star, the higher the proportion of DM it contains. This is expected because, while the NS equation of state limits the maximum mass of a NS, no such restriction applies to a DM star. Depending on the mass of the DM particle and the strength of self-interaction, a DM star’s mass can grow significantly larger choosing the above two parameters appropriately. In any case, the main conclusion remains unchanged: the admixture of DM with hadronic matter can lead to the formation of stable objects within the mass-gap region.
In Fig. 3, we confirm the stability of the three selected cases (see Fig. 2), with the help of the $P_{\rm NS}^{c}-N_{\rm DM,NS}$ diagram. In particular, we display the dependence of the number of DM particles $N_{\rm DM}$ and baryons $N_{\rm NS}$ on the pressure $P_{\rm NS}^{c}$ for equilibrium configurations of equal mass: a) $M=3.12\ M_{\odot}$, b) $M=4.18\ M_{\odot}$ and c) $M=5.24\ M_{\odot}$. The solid lines indicate the stable branches, while the dashed lines are the unstable ones. The x appearing on each curve denotes the last stable configuration.
@@
However, since the radius of these stars is not easily detectable due to the formation of the DM halo, we focus on the tidal deformability of the star. Thus, in Fig. 5, we plot the dimensional $\Lambda$ for three different masses. Obviously, the values it takes are significantly larger compared to those of NSs. Since $\Lambda$ is a quantity that can be determined from observations, it serves as an important tool for detecting these objects. It is important to note that, for a given mass, the range of $\Lambda$ values can span 2 to 3 orders of magnitude. This outcome arises from the fact that these objects, in their stable state, can vary significantly in size, particularly in terms of their radius.
@@
Moreover, we try to shed light on two critical issues concerning the creation and detection of these objects. How and when could supermassive compact objects composed of exotic fermions have formed? A relevant discussion is provided in Ref. [92]. One possibility is the accretion of DM into NSs via the primordial formation of DM clumps surrounded by ordinary matter. Another possibility involves the formation of dark compact objects resulting from DM perturbations that grow from primordial over-densities [103].
Finally, there are other possibilities for creating DM-admixed NSs with a significantly larger fraction of DM, which could accumulate gradually through capture over the NSs lifetime. One intriguing possibility is the substantial production and capture of DM during the core-collapse supernova of the neutron stars progenitor [45, 102]. In any case, there is no solid theoretical prediction that definitively rules out the formation of these hybrid NS-DM objects, where DM makes up a significantly larger proportion compared to NS matter.
The last proposition leads to the most critical question: how could we identify and detect these objects? In this case, the most powerful tool is gravitational lensing, which takes advantage of the space-time distortion caused by these objects. Another possibility is detecting a potential merger event using the well-known detectors of the LIGO-Virgo-KAGRA Collaboration. This includes two scenarios: a) the merger of two dark compact objects, or b) the merger of one dark compact object with another compact object, such as a NS or a BH. The corresponding GW signals will convey valuable information about the bulk properties of these objects-structures, including the average tidal deformability, individual masses, and chirp masses, offering insights into their structure and possibly some information about the distribution of DM in these objects [104].
Another potential detection scenario, as suggested in Ref. [99], involves a dark star in the second branch that has accreted a small amount of NS matter. Additionally, the star may have lost part of its bosonic matter due to accretion onto a second, more compact object. This process could push the star into the first unstable branch, ultimately activating the dispersion mechanism [99]. Moreover, an extensive related discussion is given in Ref. [91]. In this study, the authors suggested several ways to detect the presence of DM within NSs and provided a list of the unique signatures associated with various scenarios.
4 Concluding Remarks
@@
Summarizing this study, we propose an alternative way to explain the possible existence of compact objects in the mass-gap region. In particular, we consider that there may be hybrid states of coexisting DM and hadronic matter in the form of stable compact stars in the universe, which, among other things, cover the mass-gap region. To the best of our knowledge, this is the first attempt to explain the mass-gap using this particular hypothesis. The formation of these objects allows for multiple interpretations, and despite its significance, it lies outside the scope of this study.
Furthermore, we propose potential methods for observationally identifying these objects. While it is clear that more advanced detection techniques than those currently available are needed, we are confident that future advancements in this field will be capable of evaluating the validity of our hypothesis.
Acknowledgments
@@
This work was supported by the Czech Science Foundation (GACR Contract No. 21-24281S).